a point charge q=1C and mass 1 kg is projected with speed u=10m/s in the perpendicular direction of uniform electric field E=100 V/s.(take electric field along the direction of x- axis). the equation of trajectory of the particle is:

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

let initial velocity be along positive y-axis and point of projection be origin. y= ut, x=(1/2)(qE/m)t2, t=y/u,

thus x=(1/2)(qE/m)(y/u)2, y2= ((2m/qE)u2)x